OpenStudy (anonymous):

Is 36^-2 equal to 36^1/2?

OpenStudy (lgbasallote):

nope :D \(\huge 36^{\frac{1}{2}} = 36^{2^{-1}}\) haha see the difference? :D

OpenStudy (anonymous):

no, 36^-2=1/(36^2) and 36^1/2 =square root of 36

OpenStudy (anonymous):

Oh i see. thank you everyone :)